A road resurfacing project costs $250,000, lasts 5 years and saves $150,000 annually in patching costs. MARR is 12%. Use BCR method (2 decimal places).
Answer #1

Initial Cost = $250,000

Life = 5 years

Annual Savings = $150,000

MARR = 12%

Calculate the Benefit Cost Ratio

Using Present Worth Method

B/C Ratio = PW of Benefits ÷ PW of Cost

B/C Ratio = $150,000 (P/A, 12%, 5) ÷ $250,000

B/C Ratio = $150,000 (3.60478) ÷ $250,000

B/C Ratio = $540,717 ÷ $250,000

B/C Ratio = 2.16

The benefit cost ratio is 2.16. B.C Ratio is greater than 1, so select the project.